# Linear-limit aging times of three monoalcohols

### v1

@@ III Results and discussion
Figure 1: Equilibrium dielectric spectra at different temperatures of 2E1B, 5M2H, and 1P1P. The figure shows the imaginary part of the frequency-dependent dielectric constant $\varepsilon^{\prime\prime}$ as a function of the frequency $\nu$ plotted in a log-log plot. The insets show the molecular structures.
@@
Figure 2: Examples of dielectric loss peaks of 2E1B (green), 5M2H (orange), and 1P1P (blue). The Debye peak is marked by the letter D and the $\alpha$ peak by an $\alpha$. For 2E1B the latter is clearly visible, while it for 5M2H is manifested as a change of slope. In contrast, the 1P1P loss peak is asymmetric and looks much like those of typical non-monoalcohol organic glass formers [52, 53]. Even for 1P1P, however, have supplementary dynamic light scattering measurements revealed the existence of separate Debye and $\alpha$ peaks, albeit close to each other [38, 54].
@@
Figure 1 shows equilibrium dielectric loss spectra of the three monoalcohols for a range of temperatures. To compare the liquids we plot in Fig. 2 three representative loss spectra. 2E1B and 5M2H have large, symmetric, narrow loss peaks. These are the noted Debye loss peaks (marked by D), the above-discussed characteristic feature of monoalcohols [38]. For 2E1B there is clearly an additional high-frequency process; this is identified as the $\alpha$ process [38]. An underlying $\alpha$ process is also visible in the 5M2H data, though it is much less pronounced and merely visible as a change of slope. The 1P1P loss peak, on the other hand, has no clearly visible Debye process. Its peak is broader and asymmetric to the high-frequency side with a high-frequency slope that is not far from -0.5 [52, 53]; thus the 1P1P loss peak looks like typical peaks of van der Waals bonded liquids [52, 26]. If there is a dielectric Debye process in 1P1P, it has merged completely with the $\alpha$ process. However, a closer analysis involving also light-scattering data that are known to mainly probe the self part of the dipole time-autocorrelation function [53], reveals that the dielectric spectrum of 1P1P may also be perceived as a sum of a Debye and an $\alpha$ peak [38, 54].
The aging data and their subsequent analysis are presented in Fig. 3. The top three panels show all temperature jump data. Data are the imaginary part of the dielectric permittivity measured at a fixed frequency (1kHz for 5M2H and 10kHz for 2E1B and 1P1P), plotted as a function of the logarithm of the time after the jump was initiated. For 2E1B there are jumps to two final temperatures, for 5M2H to five final temperatures, and for 1P1P to three final temperatures. The jump magnitudes vary from 0.1 K to 2.9 K (Table 2).
Using $X=\varepsilon^{\prime\prime}(\nu_{0},t)$, where $\nu_{0}=1$ kHz or 10 kHz, as the monitored variable, the second row of Fig. 3 shows all the corresponding normalized relaxation functions $R(t)$, defined by Eq. (1), plotted as functions of time. Note that the normalized relaxation functions do not start in unity at short times; this is because there is a rapid change of the dielectric loss, which is too fast to be probed by the setup. The functions $R(t)$ do not collapse, even for jumps to the same temperature (same color). This illustrates the above-mentioned strong nonlinearity of physical aging [7]. In order to arrive at the linear-limit normalized relaxation function, $R_{\rm lin}(t)$, we employ single-parameter aging (SPA) as detailed in Sec. II.2. The fitting procedure involves just a single free parameter for each liquid, the “nonlinearity parameter” $c$ in Eq. (2). SPA uniquely determines $R_{\rm lin}(t)$ from the normalized relaxation function of any jump, $R(t)$. For each liquid, $c$ is found by searching for the best overall collapse of all calculated linear-limit aging curves. This results in $c=1.72\,\textrm{K}^{-1}$ for 2E1B and $c=1.19\,\textrm{K}^{-1}$ for both 5M2H and 1P1P.
@@
Figure 3: Aging data and SPA analysis. The left panels show data for 2E1B, the middle panels for 5M2H, the right panels for 1P1P. (a)-(c) show data for the imaginary part of the dielectric susceptibility measured at the indicated frequency after a temperature jump at $t=0$. (d)-(f) show the normalized relaxation functions corresponding to the data of (a)-(c) defined by Eq. (1) using as monitored property, $X$, the imaginary dielectric susceptibility. (g)-(i) show for each final temperature collapse of the different jump data when these are transformed into a linear aging relaxation function by means of Eq. (4). The transformation involves a single substance-specific parameter, $c$, which for each liquid is determined by best data collapse. (j)-(l) collapse the linear aging relaxation function for different target temperatures to a single one by empirically rescaling the time axis. The data collapse seen here confirms time-temperature superposition, which is a prerequisite of the TN formalism and therefore also for SPA.
@@
Table 2: List of all temperature jumps performed, starting from the initial temperature Ti and going to the final (annealing) temperature Ta. The jump magnitudes $\Delta T\equiv T_{a}-T_{i}$ are also listed.
@@
The obtained linear-limit normalized relaxation functions, $R_{\rm lin}(t)$, are shown in the third row of Fig. 3. There is a good data collapse, which validates SPA. In the fourth row we empirically scale the time axis, which for all three liquids results in a good overall data collapse. This provides a further consistency check of the analysis, because the TN formalism – and thus SPA – assumes time-temperature superposition for the linear normalized relaxation functions.
@@
Figure 4: Characteristic times $\tau$ plotted as functions of temperature for the three monoalcohols, including here also light-scattering data. The full symbols are defined as follows (compare the lower left corners that give the corresponding symbols): $\tau_{R}$ is the linear-limit aging time calculated as one over the loss-peak angular frequency of the Laplace transform of $R_{\rm lin}(t)$ determined from the nonlinear aging data by SPA via Eq. (2) (though hardly visible, the figure reports one point for each temperature jump); $\tau^{max}_{\varepsilon}$ is the inverse dielectric loss-peak angular frequency; $\tau^{\alpha}_{\varepsilon}$ is the dielectric $\alpha$ relaxation time estimated as the inverse dielectric loss-peak angular frequency (by visual inspection); the open symbols are data from the literature [55, 56, 57, 58, 56]. For 2E1B and 5M2H the linear aging times follow the $\alpha$ times, not the Debye times, in their values and temperature dependencies. For 1P1P, which is characterized by an almost complete merging of the dielectric Debye and $\alpha$ processes (Fig. 2), it would be premature to conclude although the data point to the linear-limit aging times following the Debye times better than the $\alpha$ times.
@@
From the functions $R_{\rm lin}(t)$ we extract the linear-limit aging relaxation time at each final temperature as the inverse of the loss-peak angular frequency of the Laplace transform of $R_{\rm lin}(t)$. Having done so, we can now address this paper’s scientific question: Does the temperature dependence of the linear aging time, $\tau_{R}$, follow that of the $\alpha$ process – as most other physical processes do [38, 43, 44] – or does it follow the slower Debye process [59, 60, 61]? To answer this we plot for each liquid in Fig. 4 as functions of temperature: the Debye relaxation times [55, 56, 57, 58], the $\alpha$ process relaxation times extracted by dielectric spectroscopy [55, 56, 57, 58], and the dynamic-light-scattering relaxation times [56, 57, 58].
@@
In general, different linear-response functions may have different characteristic relaxation times [41, 42, 43, 44], i.e., the same underlying process can have different relaxation times when observed with different techniques. A trivial example of this is the comparison of the compliance and modulus representation of the same data: depending on the relaxation strength, this can lead to very different relaxation times. For several liquids, a time-scale ordering has been established, where the thermal response functions, i.e., response functions where the input is a temperature modulation (dynamic heat capacity, thermal expansion coefficient, etc) – are slower than the dielectric response functions by up to one decade [43, 44]. This means that it is not possible to reach a conclusion by comparing time scales alone (although they usually do not differ by more than a decade). Thus in order to determine whether physical aging follows the Debye or the $\alpha$ process, we must not only take into account the proximity of the relaxation times but also how the corresponding relaxation times vary with temperature.
@@
The aging data cover a smaller temperature range than the other data of Fig. 4. The figure nevertheless illuminates the physics of aging and its relation to dielectric processes. For the two monoalcohols with the most pronounced and clearly separated Debye processes, 2E1B and 5M2H, the conclusion is clear: Physical aging follows the $\alpha$ process, not the Debye process. For 2E1B there is no overlap between the $\alpha$ relaxation times extracted from the dielectric spectra and the aging relaxation time, but the separation between the Debye time scale and the aging relaxation time clearly exceeds 2-3 decades and the aging time follows the temperature dependence of the dielectric $\alpha$relaxation time rather than of the dielectric Debye relaxation time. In the case of 5M2H, the $\alpha$ relaxation time found by dielectrics and light scattering is identical to the linear aging relaxation time: In this case both time scale and temperature dependence match. For 1P1P the picture is less clear. Here a separate $\alpha$ and Debye processes have been determined from data for non-linear dielectric spectroscopy [62] and light scattering [63]. While these data appear to indicate that aging follows the Debye and not the $\alpha$ process, we hesitate to conclude this firmly, because the $\alpha$ and the Debye dielectric processes of 1P1P are so close that the time scale separation in itself is inconclusive. In addition, the processes are merged to such an extent that they cannot be distinguished visually at any temperature, which also means that the temperature dependence of the two processes for all practical purposes is identical.
@@
The fact that the Debye process in 2E1B and 5M2H appears to have no relation to the linear aging timescale, while the situation is different for 1P1P, suggests a different underlying origin of the cross-correlations responsible for the Debye process. In other words, our results point to the possible existence of different classes of Debye processes: In some systems the linear aging time scale clearly follows the dielectric $\alpha$ relaxation, in other systems it is closer the dielectric Debye process. Such a difference makes sense in view of the fact that hydrogen structures evolve very differently in different systems in the form of chains [64], rings [64, 65], or in a branched [66] form. Thus 2E1B and 2E1H are classical chain-like alcohols, while 1P1P is believed to from a mixture of ring- and chain-like structures. Note also that additional cross-correlations can form in phenyl alcohols due to the pi-pi electron systems [67, 68], as has also been observed in recent simulations [69].
Only few other investigations have compared aging times to other relaxation times. Gainaru et al. found indications that the dielectric aging of 2-ethyl-1-hexanol (2E1H) and 2-ethyl-1-hexyl bromide follows the $\alpha$ process [59], while in the non-hydrogen-bonding liquid tri-butyl-phosphate (TBP) with an alcohol-like spectrum created by dielectric cross-correlation, the linear aging time follows the Debye-like mode associated with dielectric cross-correlations [60, 61].
@@ IV Summary
We have presented data for the physical aging of three monoalcohols obtained by monitoring the high-frequency dielectric loss after up and down temperature jumps of small and moderate magnitude. The aging of all three liquids conforms to SPA, which allows one to extract the linear-limit normalized aging relaxation functions $R_{\rm lin}(t)$ at each final temperature. From these functions, we determined the linear-limit aging relaxation times as the inverse of the loss-peak angular frequencies of the Laplace transform of $R_{\rm lin}(t)$. Comparing to the temperature dependence of the Debye and $\alpha$ relaxation inverse loss peak frequencies, the analysis showed that for the two liquids with best separation of the Debye and the $\alpha$ processes, 2E1B and 5M2H, physical aging follows the $\alpha$ process and not the Debye process. The third liquid, 1P1P, appears to follow the Debye process in its aging-time temperature dependence, but no definite conclusion can be made at this point because the Debye and $\alpha$ relaxation times are quite close for this system.

### v2

@@ III Results and discussion
Figure 1: Equilibrium dielectric spectra at different temperatures of 2E1B, 5M2H, and 1P1P. The figure shows log-log plots of the imaginary part of the frequency-dependent dielectric constant, $\varepsilon^{\prime\prime}$, as a function of the frequency $\nu$. The insets show the molecular structures of each monoalcohol.
@@
Figure 2: Examples of dielectric loss peaks of 2E1B (green), 5M2H (orange), and 1P1P (blue). The Debye peak is marked by the letter D and the $\alpha$ peak by an $\alpha$. For 2E1B the latter is clearly visible, while it for 5M2H is manifested as a change of slope. In contrast, the 1P1P loss peak is asymmetric and looks much like those of typical non-monoalcohol organic glass formers [57, 58]. For 1P1P supplementary dynamic light scattering measurements have revealed the existence of separate Debye and $\alpha$ peaks, albeit close to each other [38, 59].
@@
Figure 1 shows equilibrium dielectric loss spectra of the three monoalcohols for a range of temperatures. To compare the liquids we plot in Fig. 2 three representative loss spectra. 2E1B and 5M2H have large, symmetric, narrow loss peaks. These are the noted Debye loss peaks (marked by D), the above-discussed characteristic feature of monoalcohols [38]. For 2E1B there is clearly an additional high-frequency process, which is identified as the $\alpha$ process [38]. An underlying $\alpha$ process is also visible in the 5M2H data, though it is here much less pronounced and merely visible as a change of slope. The 1P1P loss peak, on the other hand, has no clearly visible Debye process. Its peak is broader and asymmetric to the high-frequency side with a high-frequency slope that is not far from -0.5 [57, 58]; thus the 1P1P loss peak looks like typical peaks of van der Waals bonded glass-forming liquids [57, 26]. If there is a dielectric Debye process in 1P1P, it has merged completely with the $\alpha$ process. However, a closer analysis involving also light-scattering data, which are known to mainly probe the self part of the dipole time-autocorrelation function [58], reveals that the dielectric spectrum of 1P1P may also be perceived as a sum of a Debye and an $\alpha$ peak [38, 59].
@@
Figure 3: Aging data and SPA analysis. The left panels show data for 2E1B, the middle panels for 5M2H, the right panels for 1P1P. (a)-(c) show data for the imaginary part of the dielectric susceptibility measured at the indicated frequency after a temperature jump at $t=0$. (d)-(f) show the normalized relaxation functions corresponding to the data of (a)-(c) defined by Eq. (1). (g)-(i) show for each final temperature collapse of the different jump data when these are transformed into a linear aging relaxation function by means of Eq. (4). This transformation involves the material-specific parameter, $c$, which for each liquid is determined to optimize the data collapse for jumps to the same temperature. (j)-(l) collapse the linear aging relaxation function for different target temperatures onto a single one by empirically rescaling the time axis. The data collapse seen here confirms time-temperature superposition, a prerequisite of the TN formalism and therefore also for SPA.
@@
Table 2: List of all temperature jumps performed, from the initial temperature Ti to the final temperature Ta. The jump magnitudes $\Delta T\equiv T_{a}-T_{i}$ are also listed.
The aging data and their analysis are presented in Fig. 3 devoting one column to each liquid. The top three panels show all temperature jump data, i.e., the imaginary parts of the dielectric permittivity measured at a fixed frequency in the kHz region plotted as a function of the logarithm of the time since the jump was initiated. For 2E1B there are jumps to two final temperatures, for 5M2H to five final temperatures, and for 1P1P to three final temperatures. The jump magnitudes vary from 0.1 K to 2.9 K (Table 2).
Using $X=\varepsilon^{\prime\prime}(\nu_{0},t)$ in which $\nu_{0}=1$ kHz or 10 kHz as the monitored variable, the second row of Fig. 3 shows all the normalized relaxation functions $R(t)$, defined by Eq. (1), plotted as functions of time. Note that the normalized relaxation functions do not start in unity at short times; this is because there is a rapid change of the dielectric loss that is too fast to be probed by the setup. This rapid change is notably larger for 2E1B than for the two other liquids. We interpret the rapid jump as reflecting the existence of a fast structural relaxation process (an aging “$\beta$ process”), which apparently is largest for 2E1B. We have no good explanation for this difference, which it could be interesting to investigate further, e.g., by comparing to the magnitude of possible dielectric $\beta$ processes.
The functions $R(t)$ do not collapse, even for jumps to the same temperature (same color), which illustrates the above-mentioned strong nonlinearity of physical aging [7]. In order to arrive at the linear-limit normalized relaxation function, $R_{\rm lin}(t)$, we employ single-parameter aging (SPA) as detailed in Sec. II.2. The fitting procedure involves one free parameter for each liquid, the “nonlinearity parameter” $c$ in Eq. (2). SPA uniquely determines $R_{\rm lin}(t)$ from the normalized relaxation function of any jump, $R(t)$. For each liquid, $c$ is found by searching for the best collapse of the calculated linear-limit aging relaxation functions for jumps to the same temperature. This results in $c=1.72\,\textrm{K}^{-1}$ for 2E1B and $c=1.19\,\textrm{K}^{-1}$ for both 5M2H and 1P1P.
@@
For these $c$ values the linear-limit normalized relaxation functions, $R_{\rm lin}(t)$, are shown in the third row of Fig. 3. There is a good data collapse, which validates SPA. In the fourth row we empirically scale the time axis, which for all three liquids results in a good overall data collapse. This provides a consistency check of the analysis, because the TN formalism – and thus SPA – assumes time-temperature superposition for the linear normalized relaxation functions.
@@
Figure 4: Characteristic times $\tau$ plotted as functions of temperature for the three monoalcohols, including here also light-scattering data. The full symbols are defined as follows (compare the lower left corners): $\tau_{R}$ is the linear-limit aging time calculated as one over the loss-peak angular frequency of the Laplace transform of $R_{\rm lin}(t)$ determined from the nonlinear aging data by SPA via Eq. (2) (though only visible upon magnification, the figure reports one point for each temperature jump); $\tau^{max}_{\varepsilon}$ is the inverse dielectric loss-peak angular frequency; $\tau^{\alpha}_{\varepsilon}$ is the dielectric $\alpha$ relaxation time estimated as the inverse dielectric loss-peak angular frequency (identified by visual inspection); the open symbols are data from the literature [60, 61, 62, 63, 61]. For 2E1B and 5M2H the linear aging times follow the $\alpha$ times, not the Debye times, in their values and temperature dependencies. For 1P1P, which is characterized by an almost complete merging of the dielectric Debye and $\alpha$ processes (Fig. 2), it would be premature to conclude anything about the temperature dependence. We note, however, that the linear-limit aging times are closer to the Debye times than to the $\alpha$ times.
@@
From the functions $R_{\rm lin}(t)$ we extract the linear-limit aging relaxation time at each final temperature as the inverse of the loss-peak angular frequency of the Laplace transform of $R_{\rm lin}(t)$. Having done so, we can now address this paper’s scientific question: Does the temperature dependence of the linear aging time, $\tau_{R}$, follow that of the $\alpha$ process – as most other physical processes do [38, 43, 44] – or does it follow the slower Debye process [45, 46, 47]? To answer this we plot for each liquid in Fig. 4 the following quantities as functions of temperature: the Debye relaxation times [60, 61, 62, 63], the $\alpha$ process relaxation times extracted by dielectric spectroscopy [60, 61, 62, 63], and the dynamic-light-scattering relaxation times [61, 62, 63].
@@
In general, different linear-response functions may have different characteristic relaxation times [41, 42, 43, 44], i.e., the same underlying process can have different relaxation times when observed with different techniques. A trivial example of this is the comparison of the compliance and modulus representation of the same data: depending on the relaxation strength, this can lead to quite different relaxation times. For several liquids, a time-scale ordering has been established in which the thermal response functions, i.e., response functions where the input is a temperature modulation (dynamic heat capacity, thermal expansion coefficient, etc) – are generally slower than the dielectric response functions by up to one decade [43, 44]. This means that it is not possible to reach a conclusion on this paper’s scientific question by comparing time scales alone. Thus in order to determine whether physical aging follows the Debye or the $\alpha$ process, we must take into account not primarily the proximity of the relaxation times, but how the relaxation times vary with temperature.
@@
The aging data cover a smaller temperature range than the other data of Fig. 4. The figure nevertheless illuminates the physics of aging and its relation to dielectric processes. For the two monoalcohols with the most pronounced and clearly separated Debye processes, 2E1B and 5M2H, the conclusion is clear: Physical aging follows the $\alpha$ process in its temperature dependence, not the Debye process. For 2E1B there is no overlap between the $\alpha$ relaxation times extracted from the dielectric spectra and the aging relaxation time, but the separation between the Debye time scale and the aging relaxation time clearly exceeds 2-3 decades and the aging time follows the temperature dependence of the dielectric $\alpha$ relaxation time rather than of the dielectric Debye relaxation time. In the case of 5M2H, the $\alpha$ relaxation time found by dielectrics and light scattering is identical to the linear aging relaxation time: In this case both time scale and temperature dependence match. For 1P1P the picture is less clear. Here separate $\alpha$ and Debye processes have been determined from data for non-linear dielectric spectroscopy [64] and light scattering [65]. These data indicate that aging follows the Debye and not the $\alpha$ process in the relaxation times, but we hesitate to conclude that this reflects a causal relation because the $\alpha$ and the Debye dielectric processes of 1P1P are so close that the time-scale determination is associated with sizable uncertainty. In particular, the temperature dependence of the two processes cannot be reliably distinguished.
@@
The fact that the Debye process in 2E1B and 5M2H appears to have no relation to the linear aging time scale while the situation possibly is different for 1P1P, suggests that different origins of the cross-correlations responsible for the Debye process are at play. In other words, our results indicate that there different classes of Debye processes may exist: In some systems the linear aging time scale clearly follows the dielectric $\alpha$ relaxation, in other systems it is closer to the dielectric Debye process. Such a difference makes sense in view of the fact that hydrogen structures evolve very differently in different systems in the form of chains [66], rings [66, 67], or in a branched [68] form. Thus 2E1B and 2E1H are classical chain-like alcohols, while 1P1P is believed to from a mixture of ring- and chain-like structures. Note also that additional cross-correlations can form in phenyl alcohols due to the pi-pi electron systems [69, 70], as has also been observed in recent simulations [71].
@@ IV Summary
We have presented data for the physical aging of three monoalcohols obtained by monitoring the high-frequency dielectric loss after up and down temperature jumps of small and moderate magnitude. The aging of all three liquids conforms to SPA. This allows to extract the linear-limit normalized aging relaxation functions $R_{\rm lin}(t)$ at each final temperature. From these functions, we determined the linear-limit aging relaxation times as the inverse of the loss-peak angular frequencies of the Laplace transform of $R_{\rm lin}(t)$. Comparing to the temperature dependence of the Debye and $\alpha$ relaxation inverse loss peak frequencies, the analysis showed that for the two liquids with best separation of the Debye and the $\alpha$ processes, 2E1B and 5M2H, physical aging follows the $\alpha$ process and not the Debye process. The third liquid, 1P1P, appears to follow the Debye process in its aging-time temperature dependence, but in this liquid no definite conclusion can be made because the Debye and $\alpha$ relaxation times are quite close.